OpenStudy (anonymous):

at a displacement of x from mean position . PE of a particle is E1 and at a displacement of y from mean position PE is E2.if at displacement x+y, PE is E then find out the relation between E,E1, E2.

OpenStudy (anonymous):

my final ans iss E=E1 +E2 but the ans is \[\sqrt{E}=\sqrt{E1}+\sqrt{E2}\]

OpenStudy (anonymous):

my teacher said u cant take sqrt both the side as u ll solve by correct mehod u ll get xactly in the sqrt form

OpenStudy (anonymous):

E1=1/2(ky^2),E2=1/2(kx^2),E=1/2k(x+y)^2 E=1/2k(x^2+y^2+2xy) E=1/2kx^2+1/2ky^2+xyk E=E2+E1+Sqrt(2E1/k*2E2/k)K E=E1+E2+2sqrt(E1*E2) E=(Sqrt(E1)+sqrt(E2))^2 sqrt(E)=sqrt(E1)+sqrt(E2)

OpenStudy (anonymous):

 I WILL GIVE U ONE CLUE I DONT KNOW THE mathematical manipulation part but pot energy at displacement X+Y=mg(H) pot energy at displacement X=mg(h1) pot energy at displacement Y=mg(h2) when you said  it would be enogh if u proved that part